OpenStudy (anonymous):

I WILL MEDAL AND FAN The regular price for a surfboard is $150 you have a coupon for 10% off, which is the better buy: 40% off the regular price or 30% off the regular price than 10% off the sale price

OpenStudy (dogzcatz):

do you have an idea?

OpenStudy (anonymous):

I think you compare them

OpenStudy (dogzcatz):

well yes, but do you know how to solve these questions

OpenStudy (anonymous):

Not really

OpenStudy (anonymous):

from the previous questions I had I think you turn the percents into decimals

OpenStudy (dogzcatz):

well how do you find the % off

OpenStudy (anonymous):

I'm not really sure

OpenStudy (photon336):

10% off means we first do this. #1 find out what 10% of 150 is #2 then take the number we find in 1 and subtract it from 150.

OpenStudy (dogzcatz):

Well What is 150x.4

OpenStudy (photon336):

You can set up a ratio and cross multiply \[\frac{ x }{ 150 } = \frac{ 10 }{ 100 }\]

OpenStudy (anonymous):

I got 60

OpenStudy (anonymous):

I'll try cross multiplying too

OpenStudy (dogzcatz):

okay 60-150 is

OpenStudy (mathstudent55):

No, 150 - 60 is...

OpenStudy (anonymous):

90

OpenStudy (dogzcatz):

okay that is the first answer

OpenStudy (mathstudent55):

$150 with a 40% discount is $90

OpenStudy (dogzcatz):

yes

OpenStudy (dogzcatz):

Do you know how to do the next step?

OpenStudy (anonymous):

i think 150 divided by .3 then the answer divided by .1

OpenStudy (mathstudent55):

NO!!!!!!!!

OpenStudy (anonymous):

I got 50

OpenStudy (mathstudent55):

Use the same process you used to take off 40%, but do it twice. First take off 30%, then take off 10%.

OpenStudy (anonymous):

okay I got 50 again

OpenStudy (dogzcatz):

no it is not 50

OpenStudy (dogzcatz):

I get a different answer

OpenStudy (mathstudent55):

To take off 40% you did two steps: 0.4 * 150 = 60, this is the 40% discount in dollars 150 - 60 = 90, this is the discounted price

OpenStudy (mathstudent55):

Do the two steps above with 30%. Then do it again with 10% to the number you get after the 30% discount.

OpenStudy (dogzcatz):

45-150=-105 turn that into 105x.1=__ subtract 105 and you should get your answer

OpenStudy (anonymous):

240

OpenStudy (dogzcatz):

where did that come from

OpenStudy (dogzcatz):

105x.1=?

OpenStudy (dogzcatz):

then that number -105

OpenStudy (dogzcatz):

that - is minus not negitive forgot to put the space

OpenStudy (dogzcatz):

negative*

OpenStudy (anonymous):

I did .3 x 150 and got 45 than subtracted 150-45 and got 105. Than I did .1 x 150 and got 15 than subtracted 150 from 15 and got135 than i subtracted both of the numbers

OpenStudy (dogzcatz):

105x.1= x then x-105=answer

OpenStudy (anonymous):

139.5

OpenStudy (mathstudent55):

I pointed out above with the 40% discount. It's not 160 - 150, it's 150 - 60 = 90

OpenStudy (anonymous):

so is it -139.5 or 139.5

OpenStudy (dogzcatz):

okay now I'm way to confused and this is a simple question

OpenStudy (mathstudent55):

Here it's the same. 0.3 * 150 = 45 150 - 45 = 105

OpenStudy (mathstudent55):

The 30% discount is done and we get 105. Now we apply the 10% discount, 0.1 * 105 = 10.5 105 = 10.5 = 94.5

OpenStudy (anonymous):

so 94.5

OpenStudy (dogzcatz):

90 94.5 yes which is the better deal?

OpenStudy (anonymous):

The 40% deal

OpenStudy (dogzcatz):

yes that is your answer

OpenStudy (anonymous):

Thank you guys so much

OpenStudy (mathstudent55):

Now we look at all our work: A single 40% discount on $150 means a final price of $90. A discount of 30% on $150 followed by a discount of 10% gives a final price of $94.50.

OpenStudy (mathstudent55):

You're welcome.

OpenStudy (mathstudent55):

There is a shortcut you can take when you need to find a discounted price. The full price of an item is 100% of the price. 100% means the full amount. If you have a 40% discount on the full price, that means you will only pay 60% of the full price since 100% - 40% = 60%. Now calculate 60% of the price to find the discounted price. Let's use the problem above as an example. 1st part: what is $150 discounted 40%? 0.6 * $150 = $90 2nd part: what is $150 discounted 30%? 0.7 * $150 = $150 What is $105 discounted 10%? 0.9 * $105 = $94.50 Notice we get the same discounted prices we got above, $90 and $94.50

OpenStudy (dogzcatz):

lot o steps for a simple problem. No prob happy to help

OpenStudy (anonymous):

Thanks I keep this for next time

OpenStudy (mathstudent55):

Great. You're welcome.

OpenStudy (mathstudent55):

@DogzCatz Great job, but be careful with 150 - 60 versus 60 - 150. It makes a big difference.

OpenStudy (dogzcatz):

Sorry confused myself and yeah that is the difference between a negative and a positive number outcome